Consider the recurrence relation an=4 an−1−4 an−2

Find the general solution to the recurrence relation and the solution when a0=−2 and a1=3.


1
Expert's answer
2022-03-08T00:40:03-0500

Let us solve the recurrence relation "a_n=4 a_{n\u22121}\u22124 a_{n\u22122}"

The characteristic equation "k^2=4k-4" is equivalent to "(k-2)^2=0,"

and hence has the roots "k_1=k_2=2."

Therefore, the general solution if of the form:

"a_n=(An+B)2^n."


Let us find the solution when "a_0=\u22122" and "a_1=3."


It follows that "-2=a_0=B" and "3=a_1=(A+B)2=(A-2)2=2A-4."

We conclude that "A=\frac{7}2,\ B=-2," and hence the solution is

"a_n=(\frac{7}2n-2)2^n."